# Multi-Microphone Speech Emotion Recognition using the Hierarchical Token-semantic Audio Transformer Architecture

### v1

@@ 4 Experimental Study
Algorithm Setup: The original HTS-AT model was pre-trained utilizing the AudioSet dataset, which comprises over two million audio samples. Each sample is 10 seconds long and is categorized into 527 distinct sound event categories. In order to adapt the pre-trained AduioSet model to RAVDESS, IEMOCAP, and CREMA-D, we adjusted the number of output classes to 7, 4, and 6, respectively. In the three datasets, pre-processing and warm-up strategy were carried out as demonstrated in [27] by providing the HTS-AT with 64 mel-bins to compute the STFT and mel-spectrograms features with 160 hop size and 1024 window size. We modified the original AdamW optimizer of HTS-AT to the traditional Adam optimizer with a learning rate of $1e^{-3}$, and batch size of 128. We set the number of epochs to 150 for all datasets. We used cross-entropy loss as the metric. We also follow the same original HTS-AT hyperparameter settings. The dimensions of the patch are set to $4\times 4$, the patch window length is 256 frames, and the attention window size is $8\times 8$. The architectural configuration comprises four network groups, each comprising 2, 2, 6, and 2 Swin-Transformer blocks, respectively. The initial patch embed is linearly projected to a dimension of $D=96$, and correspondingly, after each transformer group, the dimension increases exponentially to $7688D=$ 768, aligning seamlessly with the principles of AST. In part of our experiments, we reduced the depth of each network group by half, thereby reducing the number of trainable parameters by half to prevent overfitting. In addition, we added an early stopping strategy with a patience of 50 for RAVDESS and 25 for IEMOCAP and CREMA-D. The overall parameters for fine-tuning the datasets were 15.7M for the RAVDESS and 28.6M for IEMOCAP and CREMA-D. Since we performed fine-tuning, our models trained in less than two hours on one A6000 RTX GPU.
@@
Figure 3: Accuracy and Confidence Interval on test sets convolved with ACE RIR Lecture Room 2 ($T_{60}=1220~{}ms$). The results of two HTS-AT fine-tuned on clean and simulated RIRs datasets compared with [12] trained on clean datasets.
Results: We applied fine-tuning to the pre-trained HTS-AT AudioSet model using single-channel clean (non-reverberant) and multi-channel simulated RIRs datasets. Our starting point referred to the clean single-channel case, which will serve as one of the baselines. We also compared our results with another work that proposed a different architecture using BiLSTM together with the Attention mechanism [12]. Table 1 depicts the process of fine-tuning the model using the clean single-channel datasets. Furthermore, to complete establishing the baseline using the RAVDESS and IEMOCAP datasets, we fine-tuned another single-channel HTS-AT model by fine-tuning on the artificially reverberated datasets with a uniformly distributed $T_{60}=200-800$ ms. Then, we compared the three models by computing the Confidence Interval.333github.com/luferrer/ConfidenceIntervals The evaluation was performed on the reverberant test set using the ACE database with only one of the microphones. Examining Fig. 3 for both IEMOCAP RAVDESS datasets, it is evident that the two HTS-AT variants clearly outperform [12]. It is also clear that for RAVDESS, the HTS-AT fine-tuned with reverberant speech outperforms the HTS-AT fine-tuned using clean data. While this is also true for IEMOCAP, the differences are less significant.
@@
We now turn to the evaluation of the multi-channel schemes. In all experiments, we tested all three datasets with the utterances convolved with the three RIRs of the ACE database. We only report the results for the remote source case to emphasize the reverberation effects. The average mel-spectrogram scheme was fine-tuned using four microphones. The Patch-Embed scheme was fine-tuned using three microphones. In both cases the RIRs were generated using the simulator, as explained above. All experimental procedures followed identical training-validation splits, maintaining consistent model configurations, sizes, and hyperparameters for each dataset.
@@
Table LABEL:ACE_results provides a comparative assessment of three model variants, all fine-tuned with reverberant speech: 1) ‘HTS-AT’ - a single-channel scheme, 2) ‘Avg mel’ - mel-spectrograms averaged across four channels, and 3) ‘Sum PE’ - the Patch-Embed scheme. It is evident that the ‘Sum PE’ approach consistently outperforms the other two variants, although not by a large margin. Statistical analysis of the SER accuracy for four of the rooms is depicted in Fig. 4. Examining the accuracy bars and the confidence intervals confirms our findings regarding the consistent, albeit modest, improvements achieved by the ’Sum PE’ scheme. Importantly, these benefits come without a considerable increase in computational complexity compared to single-channel models.
(a) IEMOCAP ACE Lecture Room 1 ($T_{60}=638~{}ms$).
(b) RAVDESS ACE Meeting Room 1 ($T_{60}=437~{}ms$).
(c) CREMA-D ACE Office 1 ($T_{60}=332~{}ms$).
(d) IEMOCAP ACE Lecture Room 2 ($T_{60}=1220~{}ms$).
Figure 4: Accuracy and Confidence Interval on test sets convolved with four rooms from the ACE database. A slight improvement of the multi-channel schemes over the single-channel baseline is demonstrated, with a marginal advantage to the ‘Sum PE’ scheme.

### v2

@@ 4.2 Algorithm Setup
We fine-tuned our model starting from the original HTS-AT model that was pre-trained utilizing the AudioSet dataset, which comprises over two million audio samples. Each sample is 10 seconds long and is categorized into 527 distinct sound event categories. This pre-training provides a strong foundation for audio feature extraction, which we then adapt to our specific SER. In order to adapt the pre-trained AduioSet model to RAVDESS, IEMOCAP, and CREMA-D, we adjusted the number of output classes to 7, 4, and 6, respectively. In the three datasets, pre-processing and warm-up strategy were carried out as demonstrated in [22] by providing the HTS-AT with 64 mel-bins to compute the STFT and mel-spectrograms features with 160 hop size and 1024 window size. We modified the original AdamW optimizer of HTS-AT to the traditional Adam optimizer with a learning rate of $1e^{-3}$, and batch size of 128. We set the number of epochs to 150 for all datasets. We used cross-entropy loss as the metric. We also follow the same original HTS-AT hyperparameter settings. The patch dimensions are set to $4\times 4$, the patch window length is 256 frames, and the attention window size is $8\times 8$. The architectural configuration comprises four network groups, each comprising 2, 2, 6, and 2 Swin-Transformer blocks, respectively. The initial patch embed is linearly projected to a dimension of $D=96$, and correspondingly, after each transformer group, the dimension increases exponentially to $7688D=$ 768, aligning seamlessly with the principles of AST. In part of our experiments, we reduced the depth of each network group by half, thereby reducing the number of trainable parameters by half to prevent overfitting. In addition, we added an early stopping strategy with a patience of 50 for RAVDESS and 25 for IEMOCAP and CREMA-D. The overall number of parameters for the fine-tuned HTS-AT was 15.7M for the RAVDESS and 28.6M for IEMOCAP and CREMA-D. Since only fine-tuning is required, we could train the models in less than two hours on an A6000 RTX GPU.
4.3 Results
We begin our experimental study with the single-microphone case. We compare the fine-tuned HTS-AT model with another model developed in our group that utilizes BiLSTM combined with an Attention mechanism [12], which has achieved very good results on the RAVDESS and IEMOCAP datasets. We applied both schemes to a non-reverberant single-microphone signal. The results are presented in Table 1.
Furthermore, to complete establishing the baseline using the RAVDESS and IEMOCAP datasets, we fine-tuned another single-channel HTS-AT model by fine-tuning on the artificially reverberated datasets with a uniformly distributed $T_{60}=200-800$ ms. Then, we compared the three models by computing the Accuracy with statistical significance measure.333github.com/luferrer/ConfidenceIntervals The evaluation was performed on the reverberant test set using the ACE database only applied to one of the microphones.
@@ 4.3 Results
Fig. 3: Accuracy and Confidence Interval on test sets convolved with ACE RIR Lecture Room 2 ($T_{60}=1220~{}ms$). The results of two HTS-AT fine-tuned on either clean or simulated RIRs datasets compared with [12] trained on the clean datasets.
Examining Fig. 3 for both the IEMOCAP and RAVDESS datasets, it is evident that the two HTS-AT variants outperform the approach in [12]. It is also clear that for RAVDESS, the HTS-AT fine-tuned with reverberant speech significantly outperforms the HTS-AT fine-tuned using clean data. While this is also true for IEMOCAP, the differences are less significant. In addition, the significant gap observed for the RAVDESS dataset compared to IEMOCAP is due to differences in the nature of the datasets. RAVDESS contains more controlled, acted emotions, while IEMOCAP is more naturalistic.
@@
We now turn to the evaluation of the multi-channel schemes. In all experiments, we tested all three datasets with the utterances convolved with the three RIRs of the ACE database. We only report the results for the remote source case to emphasize the reverberation effects. The average mel-spectrogram scheme was fine-tuned using four microphones. The Patch-Embed scheme was fine-tuned using three microphones. While we used different microphone counts (3 and 4) for fine-tuning, this was done to demonstrate the flexibility of our approach in handling varying numbers of input channels. However, we acknowledge that using the same number of microphones for all models would provide a more direct comparison. In both cases, the RIRs were generated using the simulator, as explained above. All experimental procedures followed identical training-validation splits, maintaining consistent model configurations, sizes, and hyperparameters for each dataset.
@@
Table LABEL:ACE_results provides a comparative assessment of three model variants, all fine-tuned with reverberant speech: 1) ‘HTS-AT’ - a single-channel scheme, 2) ‘Avg mel’ - mel-spectrograms averaged across four channels, and 3) ‘Sum PE’ - the Patch-Embed scheme. In the table, we report the accuracy and the associated confidence interval for all seven rooms and three databases. Examining these results indicates the consistent, albeit modest, improvements achieved by the ‘Sum PE’ approach. Embedding each channel separately before summing, the model captures subtle inter-channel differences that might be lost in averaging. This allows the model to emphasize the most informative aspects of each channel for emotion recognition, leading to improved performance. Importantly, these benefits come without a considerable increase in computational complexity compared to single-channel models.